Lancelot is 6 years older than Guinevere. If in 5 years Guinevere will turn 22, how old is Lancelot now?
Found 2 solutions by math_tutor2020, greenestamps:
Answer by math_tutor2020(3817)


L = age of Lancelot right now
L-6 = age of Guinevere right now
This is because Guinevere is 6 years younger than Lancelot

Fast forward 5 years into the future to have Guinevere's age become (L-6)+5 = L+(-6+5) = L-1
Set this equal to 22 and solve for L

L-1 = 22
L = 22+1
L = 23
Lancelot is 23 year old right now.
That means Guinevere is L-6 = 23-6 = 17 years old.
As a check, 17+5 = 22 to confirm our answer.

Answer: 23 years old

Answer by greenestamps(13200)



By the elementary nature of the problem, it seems very likely that you are not expected to use formal algebra to find the answer. Use simple logical reasoning and basic arithmetic instead.

If Guinevere turns 22 in 5 years; then she is now 22-5 = 17.
Since Lancelot is 6 years older then Guinevere, he is now 17+6 = 23.